OpenStudy (anonymous):

an art collector has 62 paintings, but can display only 10 at a time. How many arrangements of 10 can be displayed? (order does matter)

OpenStudy (anonymous):

lots \[62\times61\times 60\times 59\times 58\times 57\times 56\times 55\times 54\times 53\]

OpenStudy (anonymous):

thats what i thought just seemed a little crazy thank you satellite73

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

so whats the final answer now im coming up with 2 dif

OpenStudy (anonymous):

thats what i got when i multiplied all them together but im thinking i nwas punching your numbers in wrong

OpenStudy (anonymous):

thank you luis